OpenStudy (anonymous):

Find (f+g)(x), (f-g)(x), (f•g)(x), and (f/g)(x) for f(x)=3x^2 and g(x)=5/x

OpenStudy (anonymous):

(f+g)(x)=f(x)+g(x) ((f)(g))(x)=(f(x))(g(x)) and so on.

OpenStudy (anonymous):

I know how to do that part, I'm just having some issues figuring out the multiplication and division part of it.

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

well \[(3x^2)(\frac{5}{x})=\frac{(3)(5)x^2}{x}=15x\]